### v1

# High-contrast double Bragg interferometry via detuning control

Abstract
We propose high-contrast Mach-Zehnder atom interferometers based on double Bragg diffraction (DBD) operating under external acceleration. To mitigate differential Doppler shifts and experimental imperfections, we introduce a tri-frequency laser scheme with dynamic detuning control. We evaluate four detuning-control strategies—conventional DBD, constant detuning, linear detuning sweep (DS-DBD), and a hybrid protocol combining detuning sweep with optimal control theory (OCT)—using exact numerical simulations and a five-level S-matrix model. The OCT strategy provides the highest robustness, maintaining contrast above 95% under realistic conditions, while the DS-DBD strategy sustains contrast above 90% for well-collimated Bose-Einstein condensates. These results offer practical pathways to enhancing DBD-based interferometers for precision quantum sensing and fundamental physics tests.
I Introduction
Atom interferometry (AI) enables precision measurements of inertial and fundamental physical quantities by coherently splitting and recombining atomic wave packets along distinct paths. Applications include atomic gravimetry Peters, Chung, and Chu (1999, 2001); Hu et al. (2013); Altin et al. (2013); Poli et al. (2011); Szigeti et al. (2020); Snadden et al. (1998); McGuirk et al. (2002); Rosi et al. (2015), rotation and inertial sensing Gustavson, Bouyer, and Kasevich (1997); Stockton, Takase, and Kasevich (2011); Geiger et al. (2011); Gautier et al. (2022); d’Armagnac de Castanet et al. (2024); Stolzenberg et al. (2025); Pelluet et al. (2025), precision measurements of the fundamental constants Fixler et al. (2007); Lamporesi et al. (2008); Rosi et al. (2014); Parker et al. (2018); Morel et al. (2020) and searches for new physics beyond the Standard Model such as the detection of ultralight dark matter Geraci and Derevianko (2016); Arvanitaki et al. (2018); Stadnik and Flambaum (2016); Graham et al. (2016); Du et al. (2022). The precision of these measurements hinges on extracting the phase accumulated by atoms in quantum superposition—either of motional or internal states—via fitting the signal to an a priori theoretical model.
Double Bragg diffraction (DBD) is a large-momentum-transfer technique used in atom interferometry that couples atoms to symmetric momentum states via two counter-propagating optical lattices with orthogonal polarizations (Fig. 1, left panel), first demonstrated by the group of Rasel and his collaborators within the QUANTUS consortium Ahlers et al. (2016). Compared to single Bragg diffraction (SBD) and Raman schemes, DBD offers a doubled interferometric scale factor at a given order while operating entirely within a single internal state, thereby avoiding the additional decoherence channels inherent to Raman transitions Ozeri et al. (2005); Uys et al. (2010). Moreover, its intrinsic parity symmetry Li et al. (2024) suppresses laser phase noise and systematic effects, making it well-suited for microgravity applications Barrett et al. (2016); Becker et al. (2018); Gaaloul et al. (2022); Elliott et al. (2023) and horizontal geometries Gebbe et al. (2021); Ahlers et al. (2016). However, the multi-level nature of Bragg transitions introduces unwanted parasitic paths and diffraction phases Hartmann et al. (2020); Kirsten-Siemß et al. (2023), reducing the efficiencies of the beam-splitter (BS) and the mirror (M) operations and thus limiting the overall contrast of DBD-based interferometers Ahlers et al. (2016); Gebbe et al. (2021); Stolzenberg et al. (2025).
Figure 1: Schematics of a double Bragg atom interferometer under microgravity. Left: Experimental setup of a DBD pulse using counter-propagating optical lattices $L_{1}$ and $L_{2}$ with orthogonal polarizations $\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$. Right: Real-space atomic density evolution $|\psi(z,t)|^{2}$, normalized to the initial maximum $|\psi_{max}|^{2}=\max_{z}|\psi(z,0)|^{2}$ and shown in decibel units, for conventional (C-DBD) and optimized (OCT) Mach-Zehnder interferometers with phase shifts $\Delta\phi=0$ (left column) and $\pi$ (right column), adjusted via the interrogation time $T$. Atomic density are extracted from exact numerical simulations. Red shaded regions mark the three DBD pulses.
Interferometric contrast (or visibility) is a key factor determining the sensitivity and operational limits of atom interferometers. It directly affects the maximum achievable interrogation time $T$ and effective momentum transfer $k_{\text{eff}}$, beyond which contrast loss exceeds the measurement uncertainty. At the shot-noise limit, the acceleration sensitivity of a Mach-Zehnder (MZ) interferometer scales as $\delta a=1/(\sqrt{N}k_{\text{eff}}T^{2})$, where $N$ is the number of uncorrelated atoms Bordé (1989); Kasevich and Chu (1991); Torii et al. (2000); Kasevich and Chu (1992); Storey and Cohen-Tannoudji (1994); Schleich, Greenberger, and Rasel (2013); Bongs, Launay, and Kasevich (2006); Kritsotakis et al. (2018). While near-unity contrast can be achieved in SBD-based interferometers Torii et al. (2000), DBD-based schemes are typically more sensitive to polarization errors and momentum spread, which degrade pulse efficiency and overall interferometric contrast.
In this work, we theoretically study double Bragg diffraction atom interferometers operating under a constant external acceleration, such as gravity. In Sec. II, we propose a tri-frequency laser configuration with time-dependent detuning to compensate the differential Doppler shift and improve pulse efficiency. Four detuning-control strategies are introduced: conventional DBD, constant detuning, a linear detuning sweep, and a hybrid protocol that incorporates optimal control theory (OCT) Peirce, Dahleh, and Rabitz (1988); Riahi et al. (2016); Kovachy, Chiow, and Kasevich (2012); Louie et al. (2023); Saywell et al. (2023); Goerz, Kasevich, and Malinovsky (2023); Rodzinka et al. (2024). In Sec. III, we compare the robustness of the four strategies under realistic experimental conditions. With the OCT strategy, we successfully eliminate most parasitic populations and recover near-unity contrast compared to the conventional DBD scheme (see Fig. 1, right panel). Our results offer practical means for enhancing contrast and robustness in DBD-based interferometers, with direct relevance to precision quantum sensing applications in both ground and space-based platforms.
II Double Bragg accelerometers and detuning-control strategies
II.1 Tri-frequency double Bragg accelerometers
To measure a constant linear acceleration $\bm{g}=g\hat{z}$ along the $z$-axis, we consider a MZ atom interferometer based on DBD. In microgravity or weak-acceleration regimes, a standard dual-frequency scheme suffices to drive symmetric momentum transfer in both directions Gebbe et al. (2021); Ahlers et al. (2016). However, under strong acceleration—such as gravity—the resulting differential Doppler shift $\nu_{g}=2\,k_{L}gt$ (where $k_{L}$ is the laser wave number) breaks this symmetry, preventing simultaneous resonance with both upward and downward Bragg transitions. To overcome this limitation, we propose a tri-frequency retro-reflective setup (see Fig. 2(a)), previously demonstrated in double Raman gravimeters Malossi et al. (2010). In this configuration, one of the two input frequencies (e.g., $\omega_{b}$) is replaced by a pair of frequencies $\omega_{b}\pm\nu_{D}$, where the detuning $\nu_{D}=2k_{L}a_{L}t$ linearly compensates the Doppler shift $\nu_{g}$. Due to momentum selectivity, the accelerating atoms selectively couple to four of the six laser beams, forming a pair of resonant Bragg lattices, as highlighted by the red box in Fig. 2(a).
The one-dimensional (1D) single-particle Hamiltonian describing the tri-frequency double Bragg diffraction of an atom with mass $m$ moving under a strong constant acceleration $g$, as illustrated in Fig. 2(a-b), can be written in the laboratory frame as
$H_{lab}(t)=$ $\frac{\hat{p}^{2}}{2m}+2\hbar\Omega(t)\cos\Big{(}2k_{L}\hat{z}-\int\!\nu_{D}(t)\,dt\Big{)}\times$
$\Big{\{}\cos\big{[}\Delta\omega(t)t\big{]}+\varepsilon_{pol}\Big{\}}-mg\hat{z}$, (1)
where $g$ points in the positive $z$-direction, and $\varepsilon_{pol}=|\sigma_{\perp}\cdot\sigma_{\parallel}|$ quantifies the polarization error arising from imperfect beam polarizations. $\Omega(t)$ is an effective two-photon Rabi frequency for the light-atom interaction, which can be time-dependent. For a MZ atom interferometer with triple Gaussian pulses (see Fig. 1, right panel), $\Omega(t)$ takes the following form:
$\Omega(t)=\Omega_{BS}e^{-\frac{t^{2}}{2\tau_{BS}^{2}}}+\Omega_{M}e^{-\frac{(t-T)^{2}}{2\tau_{M}^{2}}}+\Omega_{BS}e^{-\frac{(t-2T)^{2}}{2\tau_{BS}^{2}}},$ (2)
where the three light pulses are centered at time (0, $T$, $2T$) with peak two-photon Rabi frequencies ($\Omega_{BS}$, $\Omega_{M}$, $\Omega_{BS}$) and pulse widths ($\tau_{BS}$, $\tau_{M}$, $\tau_{BS}$), respectively.
Figure 2: (a) Tri-frequency laser configuration enabling double Bragg diffraction of atoms under constant acceleration $\bm{g}=g\hat{z}$, with dynamic Doppler shift compensation. The red box highlights four beams that resonantly drive DBD as the Doppler shift increases. (b) Energy level diagram of tri-frequency DBD including Doppler and AC-Stark shifts. The upward and downward Bragg transitions are driven by frequency differences $\Delta\omega_{1,2}=\omega_{b}-\omega_{a}\pm\nu_{D}$, where $\nu_{D}$ is dynamically tuned to compensate the Doppler shift $\nu_{g}=2k_{L}gt$.
It should be noted that $\Delta\omega(t)t=4\,\omega_{\mathrm{rec}}t+\Delta(t)t\equiv\phi(t_{i})+\int_{t_{i}}^{t}[\omega_{b}(t)-\omega_{a}(t)]\,dt$ in Eq. (S1) (with $\omega_{\mathrm{rec}}\equiv\hbar k_{L}^{2}/(2m)$) denotes the physically accumulated phase difference between the blue (or purple) and red lasers shown in Fig. 2(a-b), evaluated in the COM frame of the twin Bragg lattices where $\nu_{D}=0$ (see footnote [44] in Ref. Li et al. (2024)). In this COM frame, the DBD Hamiltonian (S1) transforms to
$H_{COM}(t)=$ $\frac{\hat{p}^{2}}{2m}+2\hbar\Omega(t)\cos[2k_{L}\hat{z}]\Big{\{}\cos\big{[}\Delta\omega(t)t\big{]}+\varepsilon_{pol}\Big{\}}$
$-m(g-a_{L})\hat{z}$ (3)
via a time-dependent unitary transformation $H_{COM}(t)=UH_{lab}(t)U^{\dagger}+i\hbar\dot{U}U^{\dagger}$ (see supplementary material for details). $a_{L}=\dot{\nu}_{D}/(2k_{L})$ is the COM acceleration of the twin Bragg lattices which can be fine-tuned to approach the linear acceleration $g$ such that $g-a_{L}$ is approaching zero. From Eq. (3), one can identify an effective acceleration in the new frame as $g_{eff}\equiv g-a_{L}$. For the remainder of this paper, we denote the effective acceleration $g_{eff}$ simply as $g$, and always work in the twin-lattice COM frame unless otherwise specified.
Under the microgravity assumption in the accelerated frame, i.e., $|mgz|\ll 2\hbar\Omega_{R}$ and $|g|\ll\hbar k_{L}/(mt)$ for all relevant times $t$ and positions $z$ during the interferometer sequence, the last term in Hamiltonian (3) can be treated as a perturbation and neglected during the light-matter interaction. This allows the application of the double Bragg theory developed in Ref. Li et al. (2024). During free evolution intervals between pulses, the effective acceleration leads to both a momentum shift of the atomic wave packet and an additional propagation phase dependent on the initial momentum state $|p\rangle$ and the free evolution time $T$:
$\hat{U}(T)|p\rangle$ $=e^{-\frac{iT}{\hbar}\big{(}\frac{\hat{p}^{2}}{2m}-mg\hat{z}\big{)}}|p\rangle$
$=\exp{\left(-\frac{i}{2m\hbar}(mgT^{2}p+Tp^{2})\right)}|p+mgT\rangle$
$\equiv U(p)|p+mgT\rangle,$ (4)
where the global phase proportional to $T^{3}$ has been neglected Kritsotakis et al. (2018). Since the propagation phase only depends on the momentum state $|p\rangle$, the spatially parallel trajectories will accumulate the same phase during the free fall in between successive pulses. Every applied DBD pulse will split each trajectory into different momentum classes, and hence, resulting in a coherent superposition of phase contributions from all intermediate trajectories at the final detection port.
II.2 Detuning-control strategies for high-contrast double Bragg interferometers
We now lay out four detuning-control strategies aimed at maximizing beam-splitter and mirror pulse efficiencies for Doppler-broadened wave packets with finite momentum width. Detailed pulse parameters, detuning profiles, and analyses of single-pulse efficiency and robustness are provided in the online Supplementary Material.
II.2.1 Conventional-DBD Mach-Zehnder protocol
We begin with the Mach-Zehnder sequence based on conventional double Bragg diffraction (C-DBD), where both beam-splitter and mirror pulses operate at the standard first-order DBD resonance condition $\Delta\omega=4\,\omega_{\mathrm{rec}}$ (or $\Delta(t)=0$Giese et al. (2013); Giese (2015); Ahlers et al. (2016); Li et al. (2024). The Gaussian pulse widths are individually optimized at fixed peak Rabi frequencies to balance momentum acceptance window and suppress higher-order diffraction losses at the same time. We choose $\Omega_{BS}=2.0\,\omega_{\mathrm{rec}}$ and $\Omega_{M}=2.89\,\omega_{\mathrm{rec}}$, with corresponding optimal widths $\tau_{BS}=0.47\,\omega_{\mathrm{rec}}^{-1}$ and $\tau_{M}=0.64\,\omega_{\mathrm{rec}}^{-1}$. For an input Gaussian wave packet with $\sigma_{p}=0.05\,\hbar k_{L}$ centered at $p_{0}=0$ (for BS) and $2\,\hbar k_{L}$ (for M), the efficiencies are $\eta_{BS}=97.348\%$ and $\eta_{M}=96.426\%$, respectively (see Fig. 3). The full C-DBD sequence consists of a BS–M–BS configuration with interrogation time T between pulses and serves as our baseline for comparison.
Figure 3: Comparison of beam-splitter (BS) and mirror (M) efficiencies for four strategies: conventional DBD (C-DBD), constant-detuning DBD (CD-DBD), linear-detuning-sweep DBD (DS-DBD), and a hybrid strategy combining DS-BS with OCT-mirror (OCT). Efficiencies are evaluated for an input Gaussian wave packet with a momentum width $\sigma_{p}=0.05\,\hbar k_{L}$, centered at $p_{0}=0$ for BS and $p_{0}=2\,\hbar k_{L}$ for M.
II.2.2 Constant-detuning mitigated Mach-Zehnder protocol
The second strategy employs constant detuning mitigation, where the beam-splitter and mirror pulses of the C-DBD scheme are improved by adding a fixed detuning to compensate known polarization errors and AC-Stark shift Li et al. (2024). This yields a modified resonance condition: $\Delta\omega=4\,\omega_{\mathrm{rec}}+\Delta$ with $\Delta=\text{Const.}$. Optimal detunings are found to be $\Delta_{BS}=0.27\,\omega_{\mathrm{rec}}$ for the BS and $\Delta_{M}=0$ for the mirror, using the same Gaussian pulses as in C-DBD. Without polarization error ($\varepsilon_{pol}=0$), the BS and mirror pulse efficiencies are $\eta_{BS}=99.757\%$ and $\eta_{M}=96.426\%$, respectively, for an input Gaussian wave packet with $\sigma_{p}=0.05\,\hbar k_{L}$ centered at $p_{0}=0$ and $2\,\hbar k_{L}$ (see Fig. 3). The resulting constant-detuning DBD (CD-DBD) sequence follows a BS–M–BS layout with fixed interrogation time T. This approach is effective only when polarization errors are known and the momentum distribution is narrow, ideally using box-shaped pulses.
II.2.3 Linear-detuning-sweep mitigated Mach-Zehnder protocol
The third strategy further improves the C-DBD and CD-DBD protocols by applying time-dependent linear detuning sweeps to mitigate AC-Stark and Doppler shifts, thereby improving beam-splitter and mirror pulse efficiencies Li et al. (2024). The detuning follows $\Delta(t)/\omega_{\mathrm{rec}}=(\alpha/\tau_{BS|M})(t-t_{0})+\beta$ where $t_{0}$ denotes the center of the respective Gaussian pulse, with optimized parameters $(\alpha_{BS},\beta_{BS})=(0.37,0.315)$ and $(\alpha_{M},\beta_{M})=(0.75,-4)$, using the same Gaussian pulses as in C-DBD.
This approach addresses the momentum-dependent energy shift arising from time-varying AC-Stark shifts, Doppler effects due to finite $\sigma_{p}$ or nonzero $p_{0}$, and polarization errors $\varepsilon_{pol}$—all of which may fluctuate shot-to-shot and render the resonance condition analytically intractable. To circumvent the need for precise resonance knowledge, we adopt a detuning control strategy inspired by the principle of adiabatic passage in two-level systems Steck (2025); Landau (1932); Zener and Fowler (1932), a well-established technique in nuclear magnetic resonance Bloch and Siegert (1940); Bloch (1946); Abragam (1961); Baum, Tycko, and Pines (1985); Philpott (1987). We extend this concept to the multi-level case of double Bragg diffraction via a Magnus expansion in the quasi-Bragg regime Li et al. (2024), yielding an effective two-level model between $|0\rangle=|0\hbar k_{L}\rangle$ and $|1\rangle=(|2\hbar k_{L}\rangle+|-2\hbar k_{L}\rangle)/\sqrt{2}$ (see Fig. 4(a)).
Figure 4: (a) Effective two-level system representing first-order DBD developed in Ref. Li et al. (2024). (b) Linear detuning sweep mimicking adiabatic passage for robust population transfer. The initial and final states of an ideal double Bragg beam-splitter are given by $|0\rangle=|0\hbar k_{L}\rangle$ and $|1\rangle=(|2\hbar k_{L}\rangle+|-2\hbar k_{L}\rangle)/\sqrt{2}$, respectively.
A linear sweep then enables robust and highly efficient population transfer (illustrated in Fig. 4(b)), achieving $\eta_{BS}=99.937\%$ for a DS beam-splitter at $\varepsilon_{pol}=0$—a substantial improvement over the C-DBD and within one per mil of unity, outperforming CD-DBD (Fig. 3). In contrast, the DS mirror offers only a modest gain of about $1\%$, with an efficiency $\eta_{M}=97.465\%$, still well below unity. Alternative monotonic detuning profiles (e.g., sigmoid) were also explored but offered no significant gain. The resulting DS-DBD interferometer follows a BS–M–BS sequence with interrogation time $T$. This approach is effective when the range of polarization errors is known (e.g., $\varepsilon_{pol}<5\%$), and the uncertainties in $p_{0}$ and $\sigma_{p}$ are small and under control (e.g., for well-collimated BECs).
II.2.4 Mach-Zehnder protocol with optimal control theory
The final detuning-controlled Mach-Zehnder strategy focuses on optimizing the mirror pulse—identified as the performance bottleneck in the DS-DBD protocol—using optimal control theory (OCT), using QCTRL’s Boulder Opal package Ball et al. (2021). Unlike SBD mirror pulse, the DBD mirror pulse involves a four-photon transition $|\pm 2\hbar k_{L}\rangle\rightarrow|0\hbar k_{L}\rangle\rightarrow|\mp 2\hbar k_{L}\rangle$ making it more sensitive to finite momentum spread and nonzero COM momentum Szigeti et al. (2012); Müller et al. (2008); Chiow et al. (2011); Stolzenberg et al. (2025). As illustrated in Fig. 1, a C-DBD mirror selectively filters and only reflects the central momentum component. This sensitivity is mitigated by jointly optimizing a smooth time-dependent detuning $\Delta(t)$ and all three Gaussian pulse parameters $(\Omega_{M},\,\tau_{M},\,t_{0})$, yielding an OCT mirror pulse with $(\Omega_{M},\,\tau_{M},\,t_{0})=(2.502\,\omega_{\mathrm{rec}},\,1.829\,\omega_{\mathrm{rec}}^{-1},\,3.879\,\omega_{\mathrm{rec}}^{-1})$ and an efficiency of $\eta_{M}=99.806\%$ (see Supplementary Material for detuning details), representing a substantial improvement over the DS mirror pulse (see Fig. 3). The resulting hybrid “OCT” sequence combines this mirror with two DS beam-splitters, separated by interrogation time $T$.
We also explored a fully OCT-optimized protocol, where both beam-splitter and mirror pulses are designed via OCT. However, no significant contrast improvement is observed over the DS-BS combined with OCT-mirror scheme under typical conditions with well-collimated BECs and low polarization error. A fully OCT-based approach becomes favorable only when polarization errors exceed 5%, or when large uncertainties in $p_{0}$, $\sigma_{p}$, and $\varepsilon_{pol}$ occur simultaneously.
III Contrast of double Bragg Mach-Zehnder interferometers
For spatially unresolved interferometers, i.e., the population detection after the time-of-flight ($T_{tof}$) cannot distinguish parallel spatial trajectories, either due to a short interrogation time $T$ or because the wave packet expansion during the time-of-flight $T_{tof}$ is comparable to their spatial separation, the output ports essentially correspond to different momentum states (see Fig. 1). In the far-field limit of $T_{tof}\gg T$, this correspondence becomes exact. In this case, the full interferometer can be fully described in momentum space, with its total S-matrix given by the time-ordered product of the S-matrices for individual pulses, interleaved with diagonal unitary matrices that encode the propagation phases of each momentum state accumulated between adjacent pulses.
Figure 5: (a) Mach-Zehnder atom interferometer in momentum space implemented with three DBD pulses. Input and output ports are labeled by indices $i=1,2,3$, corresponding to momentum states $|p\rangle$, $|p+2\hbar k_{L}\rangle$, and $|p-2\hbar k_{L}\rangle$, respectively. Higher-order momentum states (e.g., $|\pm 4\hbar k_{L}\rangle$) are not shown for clarity but are included in the population calculations. (b) T-scan fringe in the $|\pm 2\hbar k_{L}\rangle$-port using the DS-DBD strategy shows 97% contrast under an effective acceleration $g=0.000357k_{L}^{-1}\omega_{\mathrm{rec}}^{2}$ with an initial momentum width $\sigma_{p}=0.05\hbar k_{L}$ and COM momentum $p_{0}=0$. (c) Contrast degradation in the CD-DBD protocol due to a large initial COM momentum $p_{0}=0.1\hbar k_{L}$ and a momentum width $\sigma_{p}=0.01\hbar k_{L}$. (d) Contrast versus momentum width $\sigma_{p}$ with vanishing COM momentum and polarization error. (e) Contrast versus initial COM momentum $p_{0}$ with a momentum width $\sigma_{p}=0.05\hbar k_{L}$ and no polarization error. (f) Contrast versus polarization error $\varepsilon_{pol}$ with an initial momentum width $\sigma_{p}=0.05\hbar k_{L}$ and vanishing COM momentum. In subplots (d–f), exact numerical results are shown as symbols, while solid curves represent predictions from the five-level S-matrix theory.
The resulting S-matrix of the double Bragg Mach–Zehnder interferometer is therefore
$S^{tot}=S^{BS}U(2T,T)S^{M}U(T,0)S^{BS},$ (5)
where the contributions from all parasitic trajectories up to $\pm 4\hbar k_{L}$ momentum transfer are considered. The S-matrix of the beam-splitter pulse $S^{BS}$ in the ordered basis $\{|p\rangle,\,|p+2\hbar k_{L}\rangle,\,|p-2\hbar k_{L}\rangle,\,|p+4\hbar k_{L}\rangle,\,|p-4\hbar k_{L}\rangle\}$ is given by a $5\times 5$ matrix $(B_{ij}(p))$ with $i,\,j=1,\dots,5$, and that of the mirror pulse $S^{M}$ is given by a matrix $(M_{ij}(p))$, both of which depend on the quasi-momentum $p$ and can only be solved numerically for a Gaussian pulse with a time-dependent detuning Li et al. (2024). The two unitary matrices in above equation, given analytically by $U(T,0)=(U_{ij}(p))$ and $U(2T,T)=(U_{ij}(p+mgT))$ with $(U_{ij}(p))=diag[U(p),U(p+2\hbar k_{L}),U(p-2\hbar k_{L}),U(p+4\hbar k_{L}),U(p-4\hbar k_{L})]$ (see Eq. (4) for $U(p)$), contain the propagation phase depending on effective acceleration $g$, interrogation time $T$ as well as the input quasi-momentum $p$. With above notations, an arbitrary S-matrix element of the full interferometer can be explicitly expressed as
$S^{tot}_{ij}(g,p,T)=\sum_{k,l=1}^{5}$ $B_{il}(p_{3})U_{ll}(p_{2})M_{lk}(p_{2})U_{kk}(p_{1})\times$
$B_{kj}(p_{1}),$ (6)
[…]
$|\psi^{out}\rangle$ $=\int\!dp\,\psi(p)S^{tot}|p\rangle,$ (7)
where the final output basis is shifted due to effective acceleration to $\{|p_{3}\rangle,|p_{3}+2\hbar k_{L}\rangle,|p_{3}-2\hbar k_{L}\rangle,|p_{3}+4\hbar k_{L}\rangle,|p_{3}-4\hbar k_{L}\rangle\}$. Therefore, the output wave functions for the three main detection ports—corresponding to the signals in the black, red, and blue output ports in Fig. 5(a)—are given by $\phi_{i}(p)=\psi(p)S^{tot}_{i1}(g,p,T)$ for i=1, 2, 3i=1,\,2,\,3. The integrated population in each detection port $i$ is then given by
$P_{i}(g,T)$ $=\int_{-\hbar k_{L}}^{\hbar k_{L}}\left|\psi(p)\,S^{tot}_{i1}(g,p,T)\right|^{2}\,dp,$ (8)
respectively. Furthermore, for double Bragg interferometers, the populations in port 2 and 3 are summed to produce a signal conjugate to that in port 1, which, under ideal BS and mirror operations, takes the form of a single sinusoidal function Abend (2017):
$P_{\pm 2\hbar k}^{ideal}(g,T)=\frac{A-\mathcal{C}\cos[4k_{L}gT^{2}]}{2},$ (9)
with an offset of $A=1$ and unity contrast ($\mathcal{C}=1$) , shown as a dashed black line in Fig. 5(b–c).
For non-ideal beam-splitter or mirror operations, the contrast is generally less than unity, and the output signal typically contains multiple Fourier components Kirsten-Siemß et al. (2023). However, the contrast can still be defined, analogous to the ideal case, as the population difference between the first two non-trivial extrema of the output signal when scanning the interrogation time $T$:
$\mathbf{C}\equiv P_{\pm 2\hbar k}(g,T_{max})-P_{\pm 2\hbar k}(g,T_{min}),$ (10)
where $T_{max}$ and $T_{min}$ denote the first non-trivial maximum and minimum of $P_{\pm 2\hbar k}(g,T)\equiv P_{2}(g,T)+P_{3}(g,T)$ at a fixed $g$ as illustrated by the contrast extraction in Fig. 5(b–c). For instance, Fig. 5(b) shows that the contrast extracted for the DS-DBD strategy for an initial input state with a momentum width of $\sigma_{p}=0.05\hbar k_{L}$ and COM momentum $p_{0}=0$ without polarization error is $\mathbf{C}=97\%$. Fig. 5(c) shows the detrimental effect of a large initial COM momentum, $p_{0}=0.1\hbar k_{L}$, on the contrast, where the contrast is reduced to $\mathbf{C}=83\%$ for a narrower momentum width of $\sigma_{p}=0.01\hbar k_{L}$ without polarization error. In both cases, the results of the five-level S-matrix theory agrees perfectly with the exact numerical solutions of the Schrödinger equation based on the second-order Suzuki-Trotter decomposition Fitzek et al. (2020); Suzuki (1990) with a momentum truncation up to $\pm 14.5\hbar k_{L}$. The single Fourier fringe model (Eq. (9)) applies only in the ideal regime of small initial COM momentum and momentum width, where pulse efficiency remains nearly unaffected.
We now turn to the main results, comparing the contrast robustness of different detuning-control strategies against three key experimental imperfections: atomic momentum width $\sigma_{p}$ [Fig. 5(d)], initial center-of-mass momentum $p_{0}$ [Fig. 5(e)], and polarization error $\varepsilon_{pol}$ [Fig. 5(f)]. Under typical experimental conditions with well-controlled polarization errors ($\varepsilon_{pol}<3\%$), we find the performance ranking to be: $\text{OCT}>\text{DS-DBD}>\text{C-DBD}\approx\text{CD-DBD}$. The relative improvements of the DS-DBD and OCT protocols over C-DBD (or CD-DBD) depend on the atomic momentum width (set by the 1D effective temperature), the COM momentum (influenced by laser alignment and pulse timing), and the polarization properties of the optical beams. For a typical atom-chip-generated BEC with a 1D effective temperature of $2\,\text{nK}$ Deppner et al. (2021), corresponding to a momentum width of approximately $\sigma_{p}=0.10\hbar k_{L}$ (for $^{87}\text{Rb}$ atoms at a wavelength of $780\,\text{nm}$), the DS-DBD and OCT strategies yield relative contrast improvements of $3.4\%$ and $12.0\%$, respectively, over the C-DBD protocol (assuming perfect polarization and zero COM momentum). For a BEC after evaporative cooling with a 1D temperature below $500\,\text{pK}$ Leanhardt et al. (2003), or a momentum width about $0.05\hbar k_{L}$, the improvements reduce to $1.3\%$ (for DS-DBD) and $3.8\%$ (for OCT). In the state-of-the-art case of delta-kick collimated BECs with an effective temperature below $40\,\text{pK}$ Deppner et al. (2021); Kovachy et al. (2015), corresponding to $\sigma_{p}=0.014\hbar k_{L}$, the improvements are $0.4\%$ and $0.5\%$, respectively. With well-controlled polarization error and initial COM momentum, the DS-DBD strategy maintains contrast above $90\%$ up to $\sigma_{p}\leq 0.097\,\hbar k_{L}$, while the OCT strategy remains above $95\%$ contrast up to $\sigma_{p}\leq 0.132\,\hbar k_{L}$.
Finally, we investigate the robustness of the contrast against fluctuations in peak lattice depth for different detuning-control strategies.
Figure 6: Contrast as a function of relative peak lattice depth fluctuations for different detuning-control strategies. Symbols with error bars represent exact numerical simulations; shaded regions denote theoretical confidence intervals from the five-level S-matrix model.
Since all protocols are tuned to individually optimized parameters—including the peak lattice depth $\Omega_{BS|M}$, which is proportional to the peak laser power—they are expected to be first-order insensitive to small fluctuations in peak lattice depth. However, beyond a certain threshold, the contrast begins to degrade significantly (see Fig. 6). To quantify this threshold, we perform exact numerical simulations and plot the contrast as a function of the relative fluctuation $\sigma_{R}=\Delta\Omega_{R}/\Omega_{R}$ in Fig. 6, where $\Omega_{R}$ denotes the optimal value of peak lattice depth for either beam-splitter or mirror pulse. Each data point corresponds to an average over 10 realizations with peak lattice depths sampled from a Gaussian distribution $\Omega_{BS|M}\mathcal{N}(1,\sigma_{R}^{2})$ and a fixed momentum width of $\sigma_{p}=0.05\hbar k_{L}$ with vanishing COM momentum ($p_{0}=0$). Theoretical confidence intervals from the five-level model are shown as shaded bands, bounded by 0 and $1.05\sigma_{R}$ deviations from the optimal value. The DS-DBD protocol starts at a high contrast of $97\%$ and remains above $95\%$ up to $\sigma_{R}=3\%$, outperforming C-DBD, which holds above $95\%$ only up to $\sigma_{R}=2.5\%$. The OCT strategy exhibits the highest robustness, maintaining over $95\%$ contrast up to $\sigma_{R}=4.5\%$, outperforming both DS-DBD and C-DBD. The CD-DBD protocol shows the highest sensitivity to lattice depth fluctuations and is therefore omitted from Fig. 6 for clarity.
IV Conclusion
In summary, we have proposed and analyzed high-contrast Mach-Zehnder atom interferometers based on double Bragg diffraction under external acceleration. We introduced a tri-frequency laser scheme to compensate for differential Doppler shifts, enabling efficient beam-splitting and mirror operations. To mitigate contrast loss from experimental imperfections, such as momentum spread, center-of-mass motion, polarization errors, and peak lattice depth fluctuations, we compared the robustness of four optimized detuning-control strategies: C-DBD, CD-DBD, DS-DBD and OCT. We found that the DS-DBD strategy maintains contrast above $90\%$ for atom-chip-generated Bose-Einstein condensates, while the OCT strategy achieves the highest contrast above $95\%$ under all realistic experimental conditions. We also developed a five-level S-matrix model that accurately reproduces the interferometer output signals compared to exact numerical solutions, accounting for imperfect beam-splitter and mirror pulses as well as diffraction phases due to parasitic paths. These strategies are easily adaptable to varying experimental constraints, and alternative detection schemes—such as near-field or mid-field readout—can be naturally incorporated into the five-level S-matrix model via restricted summation indices in Eq. (6). Our results provide practical methods for enhancing contrast in DBD-based interferometers, bringing their performance on par with two-level Raman schemes and enabling their applications in high-precision quantum sensing and fundamental physics tests.
Supplementary Material
The online Supplementary Material contains two sections. Sec. S1 provides a detailed derivation of the unitary transformation from the laboratory-frame DBD Hamiltonian (Eq. (S1)) to the twin-lattice center-of-mass frame (Eq. (3)). Sec. S2 defines the beam-splitter and mirror pulse efficiencies, evaluates their robustness under experimental imperfections, and outlines the cost function used in the OCT optimization.
acknowledgement
We thank E. M. Rasel and S. Abend for the insightful discussions and thoughtful comments about the experimental feasibility. The authors gratefully acknowledge financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) CRC 1227 274200144 (DQ-mat) within Project A05, Germany’s Excellence Strategy EXC-2123 QuantumFrontiers 390837967, and through the QuantERA 2021 co-funded Project No. 499225223 (SQUEIS). We also thank the German Space Agency (DLR) for funds provided by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) due to an enactment of the German Bundestag under Grants No. 50WM2450A (QUANTUS-VI), No. 50WM2253A (AI-Quadrat), and No. 50NA2106 (QGYRO+). R.L. acknowledges the usage of LUH’s computer cluster funded by the DFG via Project No. INST 187/742-1 FUGG. N.G. and K.H. acknowledge funding by the AGAPES project - Grant No. 530096754 within the ANR-DFG 2023 Programme.
Supplementary Material: High-contrast double Bragg interferometry via detuning control
S1 Tri-frequency DBD Hamiltonian in the twin-lattice Center-of-mass frame
Here, we provide detailed derivation of the Hamiltonian for the tri-frequency double Bragg gravimeter, as presented in Eq. (3) of the main text. We begin with the lab-frame Hamiltonian, copied from Eq. (1) of the paper:
$H_{lab}(t)=$ $\frac{\hat{p}^{2}}{2m}+2\hbar\Omega(t)\cos\Big{(}2k_{L}\hat{z}-\int\!\nu_{D}(t)\,dt\Big{)}\times$
$\Big{\{}\cos\big{[}\Delta\omega(t)t\big{]}+\varepsilon_{pol}\Big{\}}-mg\hat{z}$, (S1)
where $\nu_{D}(t)=2k_{L}a_{L}t$ is the linear Doppler detuning. For brevity, we define $C(t,\varepsilon_{pol})\equiv\cos\big{[}\Delta\omega(t)t\big{]}+\varepsilon_{pol}$ for use below. We claim that this Hamiltonian can be transformed into the center-of-mass (COM) frame of the twin Bragg lattices using the following time-dependent unitary transformation:
$U=e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}-i\frac{\hat{z}}{\hbar}ma_{L}t+i\frac{\Phi(t)}{\hbar}},$ (S2)
with a properly chosen phase $\dot{\Phi}(t)=m(\frac{1}{2}a_{L}-g)\frac{1}{2}a_{L}t^{2}$ and a constant twin-lattice acceleration $a_{L}$.
To show this, let us denote the transformed Hamiltonian in the Bragg twin-lattice COM frame as $H_{COM}(t)$, which is given by
$H_{COM}(t)$ $=UH_{lab}(t)U^{\dagger}+i\hbar\dot{U}U^{\dagger}.$ (S3)
Since the exponent of $U$ contains non-commutative observables $\hat{z}$ and $\hat{p}$ with their commutator given by $\hat{z},\,\hat{p}]=i\hbar$, we apply the Baker–Campbell–Hausdorff formula Campbell (1896, 1897); Rossmann (2002) to calculate both the adjoint action of $U$ on $H_{lab}(t)$ and the time derivative of $U$. We calculate the two parts separately:
$UH_{lab}(t)U^{\dagger}$
$=$ $\frac{1}{2m}U\hat{p}^{2}U^{\dagger}+2\hbar\Omega(t)C(t,\varepsilon_{pol})\times$
$U\cos\Big{(}2k_{L}\hat{z}-\int\!\nu_{D}(t)\,dt\Big{)}U^{\dagger}-mgU\hat{z}U^{\dagger}$
$=$ $\frac{1}{2m}\big{(}U\hat{p}U^{\dagger}\big{)}^{2}+2\hbar\Omega(t)C(t,\varepsilon_{pol})\times$
$\cos\Big{(}2k_{L}\big{(}U\hat{z}U^{\dagger}-\frac{1}{2}a_{L}t^{2}\big{)}\Big{)}-mgU\hat{z}U^{\dagger}$
[…]
$-mg\Big{(}\hat{z}+\frac{1}{2}a_{L}t^{2}\Big{)},$ (S4)
where we have used the relations $U\hat{p}U^{\dagger}=\hat{p}+ma_{L}t$ and $U\hat{z}U^{\dagger}=\hat{z}+\frac{1}{2}a_{L}t^{2}$ in the last step. The second part involving the time derivative of $U$ is given by
$i\hbar\dot{U}U^{\dagger}$
$=$ $i\hbar\frac{d}{dt}\Big{(}e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}-i\frac{\hat{z}}{\hbar}ma_{L}t+i\frac{\Phi(t)}{\hbar}}\Big{)}U^{\dagger}$ (S5)
$=$ $i\hbar\frac{d}{dt}\Big{(}e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}}e^{-i\frac{\hat{z}}{\hbar}ma_{L}t}e^{i\frac{m}{4\hbar}a_{L}^{2}t^{3}+i\frac{\Phi(t)}{\hbar}}\Big{)}U^{\dagger}$ (S6)
$=$ $-a_{L}t\hat{p}+e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}}ma_{L}\hat{z}e^{-i\frac{\hat{z}}{\hbar}ma_{L}t}e^{i\frac{m}{4\hbar}a_{L}^{2}t^{3}+i\frac{\Phi(t)}{\hbar}}U^{\dagger}$
$-\frac{3}{4}ma_{L}^{2}t^{2}-\dot{\Phi}(t)$
$=$ $-a_{L}t\hat{p}+ma_{L}\hat{z}+ma_{L}\Big{[}e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}},\hat{z}\Big{]}\times$
$e^{-i\frac{\hat{z}}{\hbar}ma_{L}t}e^{i\frac{m}{4\hbar}a_{L}^{2}t^{3}+i\frac{\Phi(t)}{\hbar}}U^{\dagger}-\frac{3}{4}ma_{L}^{2}t^{2}-\dot{\Phi}(t)$ (S7)
$=$ $-a_{L}t\hat{p}+ma_{L}\hat{z}-\frac{1}{2}ma_{L}^{2}t^{2}+\frac{1}{2}mga_{L}t^{2},$ (S8)
where the relation $e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}-i\frac{\hat{z}}{\hbar}ma_{L}t}=e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}}e^{-i\frac{\hat{z}}{\hbar}ma_{L}t}e^{i\frac{m}{4\hbar}a_{L}^{2}t^{3}}$ is applied from line (S5) to line (S6) and the relation $e^{i\frac{\hat{p}}{\hbar}\frac{1}{2}a_{L}t^{2}},\hat{z}]=\frac{1}{2}a_{L}t^{2}e^{i\frac{\hat{p}}{2\hbar}a_{L}t^{2}}$ is used from line (S7) to line (S8). Combining the two parts together, one recovers Eq. (3) in the paper:
$H_{COM}(t)=$ $\frac{\hat{p}^{2}}{2m}+2\hbar\Omega(t)\cos(2k_{L}\hat{z})C(t,\varepsilon_{pol})$
$-m(g-a_{L})\hat{z}$, (S9)
with $C(t,\varepsilon_{pol})=\cos\big{[}\Delta\omega(t)t\big{]}+\varepsilon_{pol}$.
S2 Efficiencies of beam-splitter and mirror pulses for different detuning-control strategies
Here, we provide additional details regarding the beam-splitter (BS) and mirror (M) pulse efficiencies introduced in the main text. For a given quasi-momentum $p\in[-\hbar k_{L},\hbar k_{L}]$ within the first Brillouin zone, the efficiency of a double Bragg beam-splitter pulse, $\mathcal{F}_{BS}(p)$, is defined as
$\mathcal{F}_{BS}(p)=P_{|p\rangle\rightarrow|p+2\hbar k_{L}\rangle}+P_{|p\rangle\rightarrow|p-2\hbar k_{L}\rangle},$ (S10)
where $P_{|p\rangle\rightarrow|p\pm 2\hbar k_{L}\rangle}$ denotes the transition probability from the initial state $|p\rangle$ to the final state $|p\pm 2\hbar k_{L}\rangle$, respectively. Similarly, the right and left efficiencies of a double Bragg mirror pulse, $\mathcal{F}_{M}^{\pm}(p)$, are defined as
$\mathcal{F}_{M}^{+}(p)$ $=P_{|p+2\hbar k_{L}\rangle\rightarrow|p-2\hbar k_{L}\rangle},$ (S11)
$\mathcal{F}_{M}^{-}(p)$ $=P_{|p-2\hbar k_{L}\rangle\rightarrow|p+2\hbar k_{L}\rangle}.$ (S12)
Due to the parity symmetry of the double Bragg Hamiltonian in the absence of gravity Li et al. (2024), the left and right mirror efficiencies are related by $\mathcal{F}_{M}^{-}(p)=\mathcal{F}_{M}^{+}(-p)$. Without loss of generality, we choose to characterize mirror performance using the right-going efficiency $\mathcal{F}_{M}^{+}(p)$. For an initial state $|\psi(t=0)\rangle=\int\!dp\,\psi(p)|p\rangle$ with compact support in the first Brillouin zone, the integrated beam-splitter and mirror efficiencies are defined as:
$\eta_{BS}$ $=\int_{-\hbar k_{L}}^{\hbar k_{L}}\!\!\!dp|\psi(p)|^{2}\mathcal{F}_{BS}(p),$ (S13)
$\eta_{M}$ $=\int_{-\hbar k_{L}}^{\hbar k_{L}}\!\!\!dp|\psi(p)|^{2}\mathcal{F}_{M}^{+}(p).$ (S14)
S2.1 C-DBD beam-splitter and mirror pulses
The conventional double Bragg (C-DBD) beam-splitter pulse is a temporal Gaussian with parameters $(\Omega_{BS},\tau_{BS},t_{0})=(2.0\,\omega_{\mathrm{rec}},0.47\,\omega_{\mathrm{rec}}^{-1},\,0)$ where $t_{0}$ stands for center of the pulse (fixing the carrier-envelop phase at the pulse center), and a detuning $\Delta=0$. The 2-dimensional (2D) landscape of the conventional double Bragg beam-splitter efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{BS}(p,\varepsilon_{pol})$, is shown in Fig. S1(a).
Figure S1: Performance of the C-DBD beam-splitter pulse. (a) Landscape of the C-DBD beam-splitter efficiency $\mathcal{F}_{BS}(p,\varepsilon_{pol})$. (b) Transition probabilities to different momentum states for the C-DBD beam-splitter pulse with an input momentum state $|p\rangle$ with no polarization error $\varepsilon_{pol}=0$.
In Fig. S1(b), we plot the transition probabilities to different momentum states with an input momentum state $|p\rangle$ versus quasi-momentum $p$ for a fixed polarization error $\varepsilon_{pol}=0$ predicted by the five-level theory (5-LS) developed in Ref. Li et al. (2024), and compare it with the exact numerical solutions in position space using a very narrow momentum distribution $\sigma_{p}=0.01\hbar k_{L}$ centered around $p_{0}=0$ to approximate the momentum eigenstate $|p\rangle$ (abbreviated as “Exact” in the legend). For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=0$, the BS efficiency is evaluated to be $\eta_{BS}=97.348\%$.
The conventional double Bragg mirror pulse is a temporal Gaussian with parameters $(\Omega_{BS},\tau_{BS},t_{0})=(2.89\,\omega_{\mathrm{rec}},0.64\,\omega_{\mathrm{rec}}^{-1},\,0)$, and a detuning $\Delta=0$. The 2D landscape of the conventional double Bragg mirror efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{BS}(p,\varepsilon_{pol})$, is shown in Fig. S2(a).
Figure S2: Performance of the C-DBD mirror pulse. (a) Landscape of the C-DBD mirror pulse efficiency $\mathcal{F}_{M}$ versus quasi-momentum $p$ and polarization error $\varepsilon_{pol}$. (b) Transition probabilities to different momentum states after the C-DBD mirror pulse with an input state $|p+2\hbar k_{L}\rangle$ and no polarization error $\varepsilon_{pol}=0$.
In Fig. S2(b), we plot transition probabilities to different momentum states with an input momentum state $|p+2\hbar k_{L}\rangle$ versus quasi-momentum $p$ for a fixed polarization error $\varepsilon_{pol}=0$, and compare it with the exact numerical solutions in position space. For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=2\hbar k_{L}$, the mirror efficiency is evaluated to be $\eta_{M}=96.426\%$.
S2.2 CD-DBD beam-splitter pulse
The constant-detuning mitigated DBD (CD-DBD) BS pulse is a temporal Gaussian pulse sharing the same Gaussian parameters as the C-DBD BS pulse, but with an optimized constant detuning $\Delta=0.27\omega_{\mathrm{rec}}$ for quasi-momentum $p=0$ found by a numerical scan. The 2D landscape of the CD-DBD beam-splitter efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{BS}(p,\varepsilon_{pol})$, is shown in Fig. S3(a).
Figure S3: Performance of the CD-DBD BS pulse. (a) Landscape of the CD-DBD BS pulse efficiency $\mathcal{F}_{BS}$ versus quasi-momentum $p$ and polarization error $\varepsilon_{pol}$. (b) Transition probabilities to different momentum states after the CD-DBD BS pulse with an input state $|p\rangle$ and no polarization error $\varepsilon_{pol}=0$.
In Fig. S3(b), we plot transition probabilities to different momentum states with an input momentum state $|p\rangle$ for a fixed polarization error $\varepsilon_{pol}=0$, and compare it with the exact numerical solutions. For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=0$, the BS efficiency is evaluated to be $\eta_{BS}=99.757\%$.
S2.3 DS-DBD beam-splitter and mirror pulses
The linear-detuning-sweep double Bragg (DS-DBD) BS pulse is a temporal Gaussian pulse sharing the same Gaussian parameters as the CD-DBD and C-DBD BS pulse, but with an optimized linear detuning sweep $\Delta(t)/\omega_{\mathrm{rec}}=(0.37/\tau_{BS})(t-t_{0})+0.315$ found by a numerical scan. The 2-dimensional (2D) landscape of the linear-detuning-sweep double Bragg beam-splitter efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{BS}(p,\varepsilon_{pol})$, is shown in Fig. S4(a).
Figure S4: Performance of the DS-DBD beam-splitter pulse. (a) Landscape of the DS-DBD BS pulse efficiency $\mathcal{F}_{BS}$ versus quasi-momentum $p$ and polarization error $\varepsilon_{pol}$. (b) Transition probabilities to different momentum states after the DS-DBD BS pulse with an input state $|p\rangle$ and no polarization error $\varepsilon_{pol}=0$.
In Fig. S4(b), we plot transition probabilities to different momentum states with an input momentum eigenstate $|p\rangle$ versus quasi-momentum for polarization error $\varepsilon_{pol}=0$, and compare it with the exact numerical solutions. For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=0$, the DS-DBD BS efficiency is evaluated to be $\eta_{BS}=99.937\%$.
The DS-DBD mirror pulse is a temporal Gaussian with same parameters $(\Omega_{BS},\tau_{BS},t_{0})=(2.89\,\omega_{\mathrm{rec}},0.64\,\omega_{\mathrm{rec}}^{-1},\,0)$ as the C-DBD M pulse, and a linear detuning $\Delta(t)/\omega_{\mathrm{rec}}=(0.75/\tau_{BS})(t-t_{0})-4$ found by a numerical scan. The 2-dimensional (2D) landscape of the DS-DBD beam-mirror efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{M}(p,\varepsilon_{pol})$, is shown in Fig. S5(a).
Figure S5: Performance of the DS-DBD mirror pulse. (a) Landscape of the DS-DBD mirror pulse efficiency $\mathcal{F}_{M}$ versus quasi-momentum $p$ and polarization error $\varepsilon_{pol}$. (b) Transition probabilities to different momentum states after the DS-DBD mirror pulse with an input state $|p+2\hbar k_{L}\rangle$ and no polarization error $\varepsilon_{pol}=0$.
In Fig. S5(b), we plot transition probabilities to different momentum states with an input momentum state $|p+2\hbar k_{L}\rangle$ versus quasi-momentum $p$ for a fixed polarization error $\varepsilon_{pol}=0$, and compare it with the exact numerical solutions in position space. For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=2\hbar k_{L}$, the DS-mirror efficiency is evaluated to be $\eta_{M}=97.465\%$.
S2.4 OCT-mirror pulse
We want to optimize the robustness against momentum distribution and an initial COM momentum $p_{0}$. To that goal, we allow as optimization parameters: the peak Rabi frequency of the Gaussian envelope, its width, the position of the center of the Gaussian and a smooth time dependent detuning $(\Omega_{M},\tau_{M},t_{0},\Delta(t))$.
We optimize the population transfer of a Mirror interaction by optimizing the average over a range of initial momentum $p\in P=[-0.2\,\hbar k_{L},0.2\,\hbar k_{L}]$, i.e.,
$\mathrm{Cost}_{p}=$ $\bigg{\langle}\Big{|}1-|\langle+2\hbar k_{L}+p|\psi_{-2}(t=t_{f})\rangle_{p}|^{2}\Big{|}$
$+\Big{|}1-|\langle-2\hbar k_{L}+p|\psi_{2}(t=t_{f})\rangle_{p}|^{2}\Big{|}\bigg{\rangle}_{P},$ (S15)
where $|\psi_{\pm 2}(t=t_{f})_{p}\rangle$ stands for the final state after the interaction when the initial state was $|\psi_{\pm 2}(t=0)\rangle_{p}=|\pm 2\hbar k_{L}+p\rangle$.
The OCT-mirror pulse is a temporal Gaussian characterized by the parameters $(\Omega_{M},\tau_{M},t_{0})=(2.502\,\omega_{\mathrm{rec}},1.829\,\omega_{\mathrm{rec}}^{-1},3.879\,\omega_{\mathrm{rec}}^{-1})$, with a time-dependent detuning shown in the upper panel of Fig. S6. The Gaussian envelope is truncated at $T_{\mathrm{pulse}}=10\,\tau_{M}$, and the time axis is shifted as $\tilde{t}=t+T_{\mathrm{pulse}}/2-t_{0}$. The 2D landscape of the OCT-mirror efficiency versus quasi-momentum and polarization error, i.e., $\mathcal{F}_{M}(p,\varepsilon_{pol})$, is shown in the lower panel of Fig. S6.
Figure S6: Smooth OCT-mirror pulse detuning (upper) and the 2D landscape of the OCT-mirror pulse efficiency $\mathcal{F}_{M}$ versus quasi-momentum $p$ and polarization error $\varepsilon_{pol}$ (lower).
Figure S7: Transition probabilities to different momentum states after the OCT-mirror pulse with an input state $|p+2\hbar k_{L}\rangle$ and no polarization error $\varepsilon_{pol}=0$.
In Fig. S7, we plot transition probabilities to different momentum states with an input momentum state $|p+2\hbar k_{L}\rangle$ versus quasi-momentum $p$ for a fixed polarization error $\varepsilon_{pol}=0$, and compare it with the exact numerical solutions in position space. For a Gaussian momentum distribution with a finite momentum width $\sigma_{p}=0.05\hbar k_{L}$ and centered around $p_{0}=2\hbar k_{L}$, the OCT-mirror efficiency is evaluated to be $\eta_{M}=99.806\%$.
We have plotted the time-axis of the detuning in units of $\omega_{\mathrm{rec}}$. To get an idea of the duration in the laboratory, we give an example for the pulse shown in Fig. S6. For the case of $^{87}\text{Rb}$ with a wavelength of $780.1\times 10^{-9}\,m$, the duration is about $760\,\mu s$.